OpenStudy (anonymous):

. r(x)=2x+1 s(x)=x^2+2 Find the value of s(r(1)).

OpenStudy (anonymous):

r(x) = 2x + 1 s(x) = x^2 + 2 s(r(x)) = (r(x))^2 + 2 = (2x + 1)^2 + 2

OpenStudy (anonymous):

Now substitute x = 1 into s(r(x))

OpenStudy (anonymous):

(3(3(1))?

OpenStudy (anonymous):

(2(1) + 1)^2 + 2 = 3^2 + 2 = 9 + 2 = 11